Phil Pfeifer owns a business refurbishing Army Surplus calculators. He has a contract to buy the calculators from government sources and could purchase up to 5,810 a month. His bid of $5.80 per calculator had won the contract to purchase the surplus calculators. He invested $41,600 in an automated engraving machine and started selling personalized calculators through a network of army surplus stores and VFW posts.

Pricing was a problem, however. First he had to consider that, on average, his resellers charged 52% margins and were content to sell at his recommended retail prices as long as they received their margins. Second, he thought it cost him $1.24 in labor and materials to engrave customized messages. For several months he sold an average of 1,070 calculators per month at a retail selling price of $32 per customized calculator. His wife suggested he could watch more lacrosse if he charged higher prices and sold fewer calculators. Phil raised the price to $43 and saw the number of calculators sold drop to 680.

What is the profit-maximizing retail price for Phil to charge?

CALCULATED VARIABLES:

mrp = $62.57
mwb = 2,190
slope = -35
vc = $7.04
Answer #1

Formula for Profit - Maximizing retail price = 1/2 *( Cost + MRP)

Here, Cost = VC= $7.04

MRP = $62.57

So Profit-Maximizing retail price = 1/2 * ( $7.04 + $62.57 ) = $34.805

Note = All the variables were given as calculated so they are directly considered for answers. So no calculations are provided for the same.